OpenStudy (anonymous):

HELP!!! Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = one half x y = one half x + 6 y = one half x + 9

OpenStudy (anonymous):

y - 3 = (1/2)(x + 6) Distribute 1/2. y - 3 = (1/2)x + 3 Add 3 to each side. y - 3 + 3 = (1/2)x + 3 + 3 y = (1/2)x + 6